OpenStudy (anonymous):

Find a polynomial for the perimeter and for the area when one side is b+9, one side is b, and the other two sides are not labeled. The perimeter is_____. (Simplify your answer. Do not factor.)

jimthompson5910 (jim_thompson5910):

P = 2L+2W P = 2(b+9)+2b P = 2b+18+2b P = 4b+18 So the perimeter is 4b+18